HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

NEW LINE INTENSITIES FOR THE FAR INFRARED BANDS OF THE TRANS- AND CIS-CONFORMER OF NITROUS ACID (HONO), NEW DETERMINATION OF THE TRANS-CIS CONFORMER BARRIER AND ITS IMPACT ON THE ASTROPHYSICAL DETECTION OF NITROUS

ACID IN PROTOSTELLAR CLOUDS

W. TCHANA BETNGA, \underline{F. KWABIA TCHANA} , Université Paris Cité and Univ Paris Est Creteil, CNRS, LISA, F-75013 Paris, France; A. PERRIN, Laboratoire de Météorologie Dynamique/IPSL, UMR CNRS 8539, Ecole Polytechnique, Université

Paris-Saclay, RD36, 91128 Palaiseau Cedex, France; L. MANCERON, Ligne AILES, Synchrotron SOLEIL, L’Orme des Merisiers, St-Aubin BP48, 91192

Gif-sur-Yvette Cedex, France; J. VANDER AUWERA, Spectroscopy, Quantum Chemistry and Atmospheric Remote Sensing (SQUARES), C.P. 160/09, Université

Libre de Bruxelles, B-1050 Brussels, Belgium; F. HINDLE, Laboratoire de Physico-Chimie de l’Atmosphère, Université du Littoral Côte d’Opale, Dunkerque, France; A. COUTENS, Institut de Recherche en Astrophysique et Planétologie

(IRAP), Université de Toulouse, UPS, CNRS, CNES, 9 av. du Colonel Roche, 31028 Toulouse Cedex 4, France

The first goal of this work is to improve the determination of the energy difference (ΔE_{Cis−T rans}) between the ground vibrational state of the Cis- and Trans-HONO conformers of nitrous acid. For this, high resolution spectra were recorded in the 50-200 cm^{−1} spectral region at three different temperatures, 240, 270, and 296 K. The relative line intensities for the B-type transitions of pure rotational bands of Trans-HONO and Cis-HONO achieved from our measurements were combined in least squares fit computations to those measured previously by Sironneau et al.^{1} In this way, we could improve significantly the accuracy on the HONO conformer energy difference, with a value for ΔE_{Cis−T rans} = 95.8 ± 9.2 cm^{−1} compared to SIRΔE_{Cis−T rans} = 99 ± 25 cm^{−1} in the previous study of Sironneau et al. The second goal is to generate a line list with absolute line intensities for the pure rotational bands in the far infrared region of Trans-HONO and Cis-HONO, with both Aand B-type transitions. This new line list proved to be more faithful for an improved detection of HONO in Astrophysical objects^{2}.

1Sironneau V, Flaud JM, Orphal J, Kleiner I, Chelin P. Absolute line intensities of HONO and DONO in the far-infrared and re-determination of the Energy Difference between the trans- and cisspecies of nitrous acid. J Mol Spectrosc 2010;259:100-104

2Coutens A, Ligterink NFW, Loison JC, Wakelam V, Calcutt H, Drozdovskaya MN, Jørgensen JK, Müller HSP, Van Dishoeck EF, Wampfler SF. The ALMA-PILS survey: First detection of nitrous acid (HONO) in the interstellar medium. Astronomy & Astrophysics 2019;623:L13 p-number: p086  Submitted on Mon, 12 Jun 2023 16:24:05 +0200